OpenStudy (anonymous):

can someone help me understand this one? Square the binomial (6x-5)^2

OpenStudy (anonymous):

start with \[(6x-5)(6x-5)\]and then do the four multiplications

OpenStudy (anonymous):

either that, or memorize that \[(a-b)^2=a^2-2ab+b^2\] and replace \(a\) by \(6x\) and \(b\) by 5

OpenStudy (anonymous):

6 by 5 is 30 right

OpenStudy (zepp):

By\[(a-b)^2 = a^2 -2ab+b^2\] ..satellite beats me to it >.>

OpenStudy (lgbasallote):

if you'd like to learn a method in these things you can check this out http://openstudy.com/users/lgbasallote#/updates/4f9dda4ae4b000ae9ed25338 but what satellite said about expanding and multiplying is a suggested method because it is precise and you cannot go wrong with it rather than memorizing a formula :D

OpenStudy (qqstory):

how can you not know what 5 times 6 is ? o.o

OpenStudy (precal):